Table 1: Pepita Disco Performance, 2011 (UYU in millions)

Units sold 100 million units
Revenue 200
Variable costs
Materials 30
Direct labor(manufacturing,sales) 40
Opearational costs(manufacturing,inventory,delivery) 30
Other variable costs 20
Margin on Sales 80
Fixed Costs
Marketing and advertising 10
Research and development 10
Other fixed costs(e.g head office) 40
Net margin 20

a. Reduce variable costs per unit 10% and, as a result, unit sales decreased 10%?
Answer #1

Ans:

Given for 100 million units: (in millions)

Revenue = 200

Variable costs = 120

Fixed costs = 60

Workings:

Revenue per unit = $2

($200 / $100)

Variable cost per unit = $1.2

($120 / $100)

Now, variable cost per unit reduced 10% ($1.2*10% = $0.12)

i.e., variable cost per unit = $1.08 ($1.20 - $0.12)

Unit sales decreased 10%, then units to be sold = 90 million

Particulars Per unit $ Amount $ (in millions) Workings
Revenue 2 180 (90*2 = 180)
Less: variable cost (1.08) (97.20) (90*1.08 = 97.20)
Margin on sales 0.92 82.80
Less: fixed costs 60
Net margin 22.80

If any doubts or queries please comment and clarify I'll explain ASAP.